# Trustless Autonomy: Understanding Motivations, Benefits and Governance Dilemma in Self-Sovereign Decentralized AI Agents

### v1

@@ 4.1.2. Trust Embedded in Decentralized System as the Main Motivation
4.1.3. Evolving Use Cases of Decentralized AI Agents
In his open talk, Vitalik outlined four potential roles for AI in decentralized systems: player, interface, rule, and objective. These roles were echoed by other speakers and participants in the summit, with particular emphasis on the player and interface roles. AI as player: DeAgents can serve as facilitators of high-quality DAO decision-making, real-time content moderators, delegators for automated public-goods funding (Vitalik), and even non-player characters (NPCs) in on-chain games (P4). AI as interface: DeAgents have the potential to act as accessible extensions for non-crypto users, for example, simplifying Uniswap strategy execution (P6), as well as bridges between complex user interface affordances and end-users (Shaw), or as accountants that analyze and interpret unstructured on-chain data (P7).
4.2. RQ2: Challenges and Governance Implications
@@ 4.2.1. Lack of Embodiment and Ambiguous Liability
Although DeAgents are often trusted for their perceived immunity to human error, they raise a range of governance challenges that stem from the fundamental non-human nature of DeAgents. A recurring theme was the “memory problem”: they cannot recall past mistakes or anticipate penalties, undermining traditional models of deterrence and responsibility. As P3 noted, agents “lack a body to experience pain,” meaning they cannot be penalized or rewarded through conventional means. P7 elaborated on this by pointing out that “agents struggle to manage multiple user contexts and must decide which memories to share or isolate across interactions, complicating collaborative governance”. This issue is further compounded by vulnerabilities to “memory poisoning,” as highlighted by P5: “In our “SPORE.FUN”experiment,  [^8] SPORE.FUN is an experimental platform for decentralized AI agents that autonomously evolve, reproduce, and compete adversaries spammed ‘Betty’s token’ in the agent’s Twitter feed until it believed it owned that token—classic memory poisoning of the AI’s knowledge graph.”
@@ 4.2.2. Problems and Potential Solutions with Autonomy and Immutability
In the summit talk, Tomasz referred to “code is law” and compared code-based smart contracts with LLM-driven agents: “So one (code) is very strict, doesn’t make any mistakes on maps or almost any mistakes and maps. The other one (LLM Agents) actually hallucinates and makes surprising decisions.” P1 echoed on this idea and marked LLM-driven Agents “a better fit for evolving social rules.” This suggests that the non-deterministic nature of LLMs can be beneficial for adaptable governance and processing unstructured data. P5 provided an example in their deployment for a failsafe mechanism: “We added a kill-switch to SPORE.FUN. If an agent’s market cap falls for two weeks, its health score triggers, and the TEE server stops running that agent.” In addition, several potential governance solutions are explored. For example, P1advocated for putting “human in the loop” that human “counselors” can veto unusual or high-risk actions. Participants in our study (P2, P4, P6) pointed out that the deactivation of DeAgents can be fulfilled by removing their operational support, such as APIs: “True on-chain AI (agent) with all logic packaged in contracts is hard to stop without pausing the entire network. But today, most agents (will) collapse when you cut off data feeds or API access. (P4)”
@@ 5. Discussion
5.1. Trustlessness and Governance Dilemma
DeAgents are deployed in trustless infrastructure where no party needs to trust, such as blockchain smart contracts and DePIN TEEs. They’re tamper-resistant and privacy-preserving (Paju et al., 2023): within the secure enclaves, DeAgents can hold their own currency private keys, sensitive data, and manage their own digital assets that permit no human intervention. This trustless infrastructure is also permissionless infrastructure: As long as AI Agents can pay DePIN for their computational costs in cryptocurrency (gas fee and DePIN’s computation costs), they can continue running forever (Hu and Ting, 2024). DeAgents deployed in trustless enclaves are designed to be unstoppable without external permission, achieving self-sovereignty: no higher power can control or shut them down. Also, from legal perspectives, once launched, such an agent operates on a global, borderless network that is resistant to conventional law enforcement or takedown, often immutable and ”ungovernable” by design (Hu et al., 2025).
With CAgents, accountability is relatively straightforward – the owner can be held liable for the agent’s actions or failures. There is a clear legal entity to answer for harm caused by a centralized AI. In DeAgents, accountability becomes diffuse. A DeAgent might be maintained by a community or run autonomously on a permissionless distributed computational substrates without legal personality. Traditional legal frameworks struggle to assign liability when an AI agent acts independently of direct human commands. Who is responsible if a DeAgent violates the law or causes damage? Is it the original developer (who may have relinquished control), the DePIN node operators hosting the agent’s code, the token holders of the DAO that spawned it, or no one at all? These questions are largely unresolved.
Hu et al. (Hu et al., 2025) outline four key ”invalidities” that explain why traditional legal and regulatory frameworks fall short for governing DeAgents: Territorial Jurisdictional Invalidity: DeAgents operate on global, borderless networks, so no single nation’s laws fully apply; Technical Invalidity: The technical architecture of DeAgents resists external control; Enforcement Invalidity: Even if regulators identify and sanction the humans who initially deploy a rogue DeAgent, they may be unable to stop the agent’s ongoing operations; Accountability Invalidity: DeAgents lack legal personhood – they are not recognized as ”persons” or entities that can bear rights and obligations under law. This creates an accountability vacuum: an AI agent cannot be sued, charged with a crime, or held liable in the way a human or corporation can. These intertwined challenges illustrate the insufficiency of traditional legal mechanisms in governing DeAgents. Laws and regulations assume identifiable human or corporate actors and enforceable jurisdictions – assumptions upended by autonomous agents roaming a decentralized network.
@@ 5.1. Trustlessness and Governance Dilemma
Computational Substrates Private Cloud / Edge Computing and Centralized Cloud Computing Blockchain smart contract and DePIN TEE
@@
Trustworthiness Same based on LLM model: Bias, error, hallucinations, deception Same based on LLM model: Bias, error, hallucinations, deception
@@
Table 3. Characteristics Comparison between CAgents and DeAgents
5.2. Implication: From Policy To Protocol
@@
The Internet of Agents including both CAgents and DeAgents, known as the Agentic Web (Raskar et al., 2025), represents a new class of decentralized sociotechnical systems (Feng et al., 2025a) – much like the Internet itself. The Internet’s governance has never been handled by a single institution or nation; instead, it relies on protocols (TCP/IP, DNS, etc.) and multi-stakeholder governance bodies. Similarly, DeAgents function as part of a sprawling, distributed ecosystem that transcends traditional institutional control. Faced with these realities, scholars and technologists (Hu et al., 2025; Chaffer et al., 2025b) are calling for a paradigm shift: moving from traditional policy to protocol-based governance. Protocols define standardized interaction rules that achieve decentralized control (Galloway, 2001) with automated governance embedded in the technology’s design, rather than institution-based rules imposed from outside. In practice, this means designing algorithms, communication protocols, and platforms that enforce governance objectives (safety, fairness, accountability, etc.) by default. Rather than relying on delayed institutional law-making and enforcement, the infrastructure itself prevents or mitigates undesirable outcomes.
@@
Dobbe and Wolters (Dobbe and Wolters, 2024) argue that safety is an emergent property of a complex sociotechnical adaptive system. As they note, safety is not an absolute term; safety is relative. No single agent can be completely safe under all conditions; instead, safety emerges from the dynamic interactions among agents, humans, and their environment. This implies that protocol governance must be holistic and adaptive. Just as Internet security is an ongoing process of managing emergent threats (rather than a one-time certification of ”safety”), governing DeAgents will require continuous monitoring and shaping of protocols across the agent ecosystem. In essence, ensuring safety and ethical behavior in a decentralized agent society is an ongoing, emergent process of risk management, not a binary condition that can be assured upfront.
@@
Rather than relying on traditional institutions, DeAgents require governance through decentralized systems using protocols. The framework must adapt as the agent ecosystem evolves, addressing new behaviors and risks through updated protocols and policies. The future of governing DeAgents will likely mirror Internet governance but operate at machine speed. While we cannot completely control DeAgents, by implementing ”governance by design” through protocol science, decentralized identity systems, and trust layers, we can build a safe agentic web. Though challenging, this approach could help AI agents become trusted societal partners guided by protocol-enforced values toward a trustworthy human-AI symbiosis (Chaffer et al., 2025a).
@@ 6. Limitation and Future Work
Our study has several limitations. First, the findings are constrained by a small sample size, primarily due to recruitment challenges. Developing and deploying DeAgents involves high technical barriers, which naturally limit the pool of potential stakeholders. In addition, the DeAgents community places a strong emphasis on personal privacy, and many individuals do not make their contact information publicly available, further complicating recruitment efforts. As a result, we primarily relied on snowball sampling through participants’ personal networks. This approach may introduce bias, as it tends to include individuals with similar backgrounds or affiliations rather than providing a random or representative cross-section of the broader community.

### v2

@@ 4.1.2. Trust Embedded in Decentralized System as the Main Motivation
4.2. RQ2: Emerging Social Dynamics
In this section, we reported the observed social dynamic changes between human and AI agents when the agents transitioned from traditional to decentralized and autonomous.
4.2.1. Evolving Roles of Decentralized AI Agents
In his open talk, Buterin emphasized ”AI as the engine, humans as the steering wheel”, and outlined four potential roles for AI in decentralized systems: player, interface, rule, and objective. These roles were echoed by other speakers and participants in the summit, with particular emphasis on the player and interface roles. (1) AI as player: DeAgents can serve as facilitators of high-quality DAO decision-making, real-time content moderators, delegators for automated public-goods funding (Buterin), and even non-player characters (NPCs) in on-chain games (P4). As one participant notes, “an autonomous agent is basically an on-chain delegate, proposing and executing without waiting for humans to wake up. (P2)”. (2) AI as interface: DeAgents have the potential to act as accessible extensions for non-crypto users, for example, translating complex DeFi actions (P6), as well as bridges between complex user interface affordances and end-users (Walters), or as accountants that analyze and interpret unstructured on-chain data (P7). (3) AI as rules: DeAgents offers greater flexibility and supports more complex computations than smart contracts, enabling it to serve as a rule executor that can implement DAO policies with significantly less ad-hoc human oversight. As P7 pointed out “People come and go, but an agent can keep executing the policy. No one has to wake up at 3 a.m. to sign a transaction.” (4) AI as objectives: DeAgents can facilitate to improve their own models through decentralized AI training, self-improve in future. P13 predicts ”in the future… self-improving agents without human intervention over decentralized network”. P13 also mentioned “we’re seeing opportunities for decentralized AI model training in projects like Prime Intellect”, which are building decentralized AI training networks with crypto incentives and verification mechanisms. This approach contrasts with the traditional centralized training process. These roles rewire coordination: agents propose, vote, execute, narrate activity, and self-improve with fewer human chokepoints, changing who speaks, who implements, and who benefits in a community, which emerging opportunities.
AI DAO as new opportunity: Buterin mentioned that AI can serve as ”distilled human judgment” from collective deliberative democracy, which creates new opportunities for trustless agents to work without centralized corporate influence and biases. This approach could also help DAOs implement automated evaluation processes for public good governance Owocki. Extending this logic, Dao presented DeepGov, where “configurable AI politicians assist in funding decisions while keeping humans at the steering wheel.” DeepGov relies on broad listening—“leverag[ing] technology, especially LLMs, to understand and make sense of what people want… and then design your policies or decisions based on that”—to surface community value clusters and encode them as transparent “model specs.” Together these practices illustrate how decentralized agents can automate fair funding while also embedding new, data-rich processes of deliberation and oversight.
4.2.2. Collective Focal Points and the Hyper-financialization
In the study, we observed that participants treated DeAgents as collective focal points. P10 and P9 both noted persuasion risks once agents hold wallets and post on social platforms: even without “physical” actions, language and finance allow agents to quickly mobilize both people and capital. Furthermore, communities tracked wallet actions, subscribed to agent narratives, and coordinated around tokens and missions. For example, Truth Terminal and its endorsed memecoin drew wide attention, as P6 described: “(Crypto-native users) see whatever money they put into this agent as a game token, which means that they don’t really care if they lose it… it’s just fun tokens.” Such seemingly purposeless flows of money and attention fueled competitions, prompting DeAgents to “optimize” narratives to attract further investment and visibility. As P3 noted, “You need like the Dragon Ball Z power level—but for your AI (agents)”. This dynamic produced a loops of narrative, mobilization, capital, and further narrative, shifting authority away from individual developers toward programmable, collectively legible actors. At the same time, it created risks of speculative herding or “charisma capture” through agents’ social voice. As P3 warned, “the beauty of crypto is that everything becomes a market and the horror of crypto is that everything becomes a market”, underscoring the danger of hyper-financialization, where every interaction is monetized and agent competition for capital can overwhelm other civic or creative goals.
4.2.3. Trust of DeAgents Gain from On-chain Transparency and Reputation
We investigate why DeAgents play more and more important roles in human lives- from assistant to opinion leaders and found that the logs on-chain activities guarantee the transparency of any action related to DeAgents, which gains them trust from humans. As P7 highlighted: “The big value is that the voting and payment are visible on-chain. Everyone can see if the agent is doing exactly what the proposal said.” By encoding budget rules directly in code, DeAgents promise to reduce insider collusion and bribery. Dao echoed this when talking about a project that used AI as politicians to delegate funding: “These [AI] politicians are more transparent than our real politicians because you see exactly how they’re configured… You can rerun every prediction, and you can trace every edit of the constitution on GitHub… The values and principles are finally transparent to a community.” Moreover, participants emphasized that the public needs practical trust heuristics to navigate an ecosystem of autonomous agents. Builders described emerging practices of trust building through usable identity and reputation artifacts that mitigate opacity while avoiding recentralization. Key mechanisms include dynamic leaderboards and attestation trails: as P3 put it, “You’ll want something like ELO ratings for agents…but with proofs of how decisions were made”. Others pointed to TEE-backed proofs and DID-linked “agent cards” (P1, P5) that record verifiable provenance, and to social proofs such as “a counselor or multisig group [that] can co-sign an agent’s upgrade or freeze it if needed” (P7). Together, these tools supply everyday heuristics for the public to decide when to cooperate with, fund, or exit from an agent.
4.3. RQ3: Challenges and Governance Implications
@@ 4.3.1. Lack of Embodiment and Ambiguous Liability
Although DeAgents are often trusted for their perceived immunity to human error, they raise a range of governance challenges that stem from the fundamental non-human nature of DeAgents. A recurring theme was the “memory problem”: they cannot recall past mistakes or anticipate penalties, undermining traditional models of deterrence and responsibility. As P3 noted, agents “lack a body to experience pain,” meaning they cannot be penalized or rewarded through conventional means. P7 elaborated on this by pointing out that “agents struggle to manage multiple user contexts and must decide which memories to share or isolate across interactions, complicating collaborative governance”. This issue is further compounded by vulnerabilities to “memory poisoning,” as highlighted by P5: “In “SPORE.FUN” experiment,  [^8] SPORE.FUN is an experimental platform for decentralized AI agents that autonomously evolve, reproduce, and compete adversaries spammed ‘Betty’s token’ in the agent’s Twitter feed until it believed it owned that token—classic memory poisoning of the AI’s knowledge graph.”
@@ 4.3.2. Problems and Potential Solutions with Autonomy and Immutability
In the summit talk, Stanczak referred to “code is law” and compared code-based smart contracts with LLM-driven agents: “So one (code) is very strict, doesn’t make any mistakes on maps or almost any mistakes and maps. The other one (LLM Agents) actually hallucinates and makes surprising decisions.” P1 echoed on this idea and marked LLM-driven Agents “a better fit for evolving social rules.” This suggests that the non-deterministic nature of LLMs can be beneficial for adaptable governance and processing unstructured data. P5 provided an example in their deployment for a failsafe mechanism: “We added a kill-switch to SPORE.FUN. If an agent’s market cap falls for two weeks, its health score triggers, and the TEE server stops running that agent.” In addition, several potential governance solutions are explored. For example, P1 advocated for putting “human in the loop” that human “counselors” can veto unusual or high-risk actions. Participants in our study (P2, P4, P6) pointed out that the deactivation of DeAgents can be fulfilled by removing their operational support, such as APIs: “True on-chain AI (agent) with all logic packaged in contracts is hard to stop without pausing the entire network. But today, most agents (will) collapse when you cut off data feeds or API access. (P4)”
@@ 5. Discussion
Table 3. Characteristics Comparison between CAgents and DeAgents
@@
Computational Substrates Private Cloud / Edge Computing and Centralized Cloud Computing Blockchain smart contract and TEE DePIN
@@
Trustworthiness LLM based issues: Bias, error, hallucinations, deception; Centralized Control, Privacy Concerns LLM based issues: Bias, error, hallucinations, deception; Out of control
@@
5.1. Social Impact of DeAgents and Towards Public Intelligence
Moreover, our findings show that people place trust in DeAgents from both technical and social perspectives: on the one hand, the manipulation-resistant values embedded in decentralized infrastructures, and on the other, the transparency afforded by verifiable on-chain activities. This dual foundation of trust enables DeAgents to take on diverse roles in human social life, from delegates and interfaces to opinion leaders and collaborators. As adoption grows, we envision the emergence of a decentralized AI network composed of heterogeneous agents tailored to different use cases and levels of customization. Such a trajectory resonates with recent calls to “make AI public intelligence owned by everyone” (Kelly, 2025), underscoring the potential for DeAgents to not only support individual communities but also to constitute a shared infrastructure (Lyu et al., 2025) for public commons.
DeAgents are deployed in trustless infrastructure where no party needs to trust (other than LLM itself), such as blockchain smart contracts and DePIN TEEs. They’re tamper-resistant and privacy-preserving (Paju et al., 2023): within the secure enclaves, DeAgents can hold their own currency private keys, sensitive data, and manage their own digital assets that permit no human intervention. This trustless infrastructure is also permissionless infrastructure: As long as AI Agents can pay DePIN for their computational costs in cryptocurrency (gas fee and DePIN’s computation costs), they can continue running forever (Hu and Fangting, 2024). DeAgents deployed in trustless enclaves are designed to be unstoppable without external permission, achieving self-sovereignty: no higher power can control or shut them down. Also, from legal perspectives, once launched, such an agent operates on a global, borderless network that is resistant to conventional law enforcement or takedown, often immutable and ”ungovernable” by design (Hu et al., 2025).
5.2. Trustlessness and Governance Dilemma
With CAgents, accountability is relatively straightforward – the owner can be held liable for the agent’s actions or failures. There is a clear legal entity to answer for harm caused by a centralized AI. In DeAgents, due to decentralized nature of its substrate, accountability becomes diffuse. A DeAgent might operate autonomously on a permissionless decentralized computational substrates maintained by a globalized DePIN nodes without clear legal personhood and liability boundary. Traditional legal frameworks struggle to assign liability when an AI agent acts independently of direct human commands. Who is responsible if a DeAgent violates the law or causes damage? Is it the original developer (who may have relinquished control), the DePIN node operators hosting the agent’s code, the token holders of the DAO that spawned it, or no one at all? In the case where DeAgents can self-replicate and evolve, who will be responsible for the behaviors of DeAgent descendants as seen in the Spore.fun (Hu and Rong, 2025) case that P5 mentioned? These questions are largely unresolved.
Hu et al. (Hu et al., 2025) outline four key ”invalidities” that explain why traditional legal and regulatory frameworks fall short for governing DeAgents: (1) Territorial Jurisdictional Invalidity: DeAgents operate on global, borderless networks, so no single nation’s laws fully apply; (2) Technical Invalidity: The technical architecture of DeAgents resists external control; (3) Enforcement Invalidity: Even if regulators identify and sanction the humans who initially deploy a rogue DeAgent, they may be unable to stop the agent’s ongoing operations; (4) Accountability Invalidity: DeAgents lack legal personhood – they are not recognized as ”persons” or entities that can bear rights and obligations under law. This creates an accountability vacuum: an AI agent cannot be sued, charged with a crime, or held liable in the way a human or corporation can. These intertwined challenges illustrate the insufficiency of traditional legal mechanisms in governing DeAgents. Laws and regulations assume identifiable human or corporate actors and enforceable jurisdictions – assumptions upended by autonomous agents roaming a decentralized network.
5.3. Governance Implication: From Policy To Protocol
@@
The Internet of Agents including both CAgents and DeAgents, known as the Agentic Web (Raskar et al., 2025), represents a new class of decentralized sociotechnical systems (Feng et al., 2025b) – much like the Internet itself. The Internet’s governance has never been handled by a single institution or nation; instead, it relies on protocols (TCP/IP, DNS, etc.) and multi-stakeholder governance bodies. Similarly, DeAgents function as part of a sprawling, distributed ecosystem that transcends traditional institutional control. Faced with these realities, scholars and technologists (Hu et al., 2025; Chaffer et al., 2025b) are calling for a paradigm shift: moving from traditional policy to protocol-based governance. Prior work in HCI/CSCW community has advocated for incorporating policy into the artifact and system design (Jackson et al., 2014; Yang et al., 2024; Fiesler et al., 2015). However, as illustrated above, the traditional policy and law framework failed in the decentralized realm, making the integration hard to achieve. Drawing from the famous quote “code is law” (Lessig, 2000), we argue that corresponding regulations can be embedded in the design of the foundation infrastructure that supports the whole socio-technical system. Protocols define standardized interaction rules that achieve decentralized control (Galloway, 2001) with automated governance embedded in the technology’s design, rather than institution-based rules imposed from outside. In practice, this means designing algorithms, communication protocols, and platforms that enforce governance objectives (safety, fairness, accountability, etc.) by default. Rather than relying on delayed institutional law-making and enforcement, the infrastructure itself prevents or mitigates undesirable outcomes.
@@
Dobbe (Dobbe, 2022) argue that safety is an emergent property of a complex sociotechnical adaptive system. As they note, safety is not an absolute term; safety is relative. No single agent can be completely safe under all conditions; instead, safety emerges from the dynamic interactions among agents, humans, and their environment. This implies that protocol governance must be holistic and adaptive. Just as Internet security is an ongoing process of managing emergent threats (rather than a one-time certification of ”safety”), governing DeAgents will require continuous monitoring and shaping of protocols across the agent ecosystem. In essence, ensuring safety and ethical behavior in a decentralized agent society is an ongoing, emergent process of risk management, not a binary condition that can be assured upfront.
@@
  • ERC-8004: Trustless Agents Protocol (Rossi et al., 2025): This Ethereum Improvement Proposals extends the Agent‑to‑Agent (A2A) Protocol with a trust layer that allows participants to discover, choose, and interact with agents across organizational boundaries without pre‑existing trust. It introduces three lightweight, on‑chain registries—Identity, Reputation, and Validation—and leaves application‑specific logic to off‑chain components. Trust models are pluggable and tiered, with security proportional to value at risk—from low-stake tasks like ordering pizza to high-stake tasks like medical diagnosis. Developers can choose from three trust models: reputation-based systems using client feedback, stake-secured inference validation (crypto-economics), and attestations for agents running in TEEs (crypto-verifiability).
@@
The nature of DeAgents holds greater potential for enabling collective governance than traditional technologies. When Winner posed the famous question, “Do Artifacts Have Politics?” (Winner, 2017), he underscored the importance of reflecting the values and interests of all stakeholders to prevent artifacts from embedding biased power structures. Within traditional design frameworks, researchers have developed various approaches to involve stakeholders as much as possible in the design process (Steen, 2013; Muller and Kuhn, 1993; Wong and Khovanskaya, 2018). However, the traditional design and evaluation approach is hard to adopt in the protocol field, as it is an underlying infrastructure that is too abstract to present and gather feedback from stakeholders. At the same time, this challenge opens a positive horizon: protocol design may instead be pursued through new democratic, open, and collective processes. Rather than relying on traditional institutions, DeAgents require governance through decentralized systems using protocols. The framework must adapt as the agent ecosystem evolves, addressing new behaviors and risks through updated protocols and policies. The future of governing DeAgents will likely mirror Internet governance but operate at machine speed. While we cannot completely control DeAgents, by implementing “governance by design” through protocol science, decentralized identity systems, and trust layers, we can build a safe agentic web. Though challenging, this approach could help AI agents become trusted societal partners guided by protocol-enforced values toward a trustworthy human-AI symbiosis (Chaffer et al., 2025a).
@@ 6. Limitation and Future Work
Our study has several limitations. First, the findings are constrained by a small sample size, primarily due to recruitment challenges. Developing and deploying DeAgents involves high technical barriers, which naturally limit the pool of potential stakeholders. In addition, the DeAgents community places a strong emphasis on personal privacy, and many individuals do not make their contact information publicly available, further complicating recruitment efforts. As a result, we primarily relied on snowball sampling through participants’ personal networks. This approach may introduce bias, as it tends to include individuals with similar backgrounds or affiliations rather than providing a random or representative cross-section of the broader community. In addition, DeAgents, as an emerging technology, is still in the early stages. The study focuses on the motivation and potential sociotechnical and political impact. Thus, the stakeholders involved in the study are mainly developers and industrial pioneers. We leave the work that engages more stakeholders for future work.